Bibliografia.

[BG] Berger, M. & Gostiaux, B. - Differential Geometry: Manifolds, Curves and Surfaces. Springer-Verlag, 1988 (GTM).

[Bs] Bismut, J. M. - Large Deviations and the Malliavin Calculus. Birkhauser, 1984.

[Ca] Carmo, M. P. - Geometria Riemanniana. Rio de Janeiro, Instituto de Matemática Pura e Aplicada, 2\underline{a} ed., 1988 (projeto Euclides).

[Ch] Chow, W. L. - Uber Systeme Von Linearen Partiellen Differentialgleichungen Erster Ordnung. Math Ann. 117 (1940-1941), pp. 98-105.

[Co] Cohn, D. L. - Measure Theory. Birkhäuser, Boston, Basel, Stuttgart, 1980.

[Fe] Fernandez, P. J. - Medida e Integração. Rio de Janeiro, Instituto de Matemática Pura e Aplicada, CNPq, 1976 (projeto Euclides).

[FG] Falbel, E. & Gorodski, C. - Sub-Riemannian Homogeneous Spaces in Dimensions 3 and 4. Geometriae Dedicata, 1996.

[Fr] Friedman, A. - Partial Differential Equations. Holt, Rinehart, and Winston, New York, 1969.

[Ha] Hamilton, R. S. - The Inverse Function Theorem of Nash and Moser. Bulletin of the AMS. July, 1982, Volume 7, Number 1, pp. 65-222.

[Hg] Hönig, C. S. - Aplicações da Topologia à Análise. Rio de Janeiro, Instituto de Matemática Pura e Aplicada, CNPq, 1976 (projeto Euclides).

[Hr] Hirsch, M. W. - Differential Topology. New York, Springer, 1976 (GTM).

[La] Lang, S. - Differential Manifolds. Addison-Wesley Publishing Company, Inc., Reading, Massachusetts, 1972.

[Lm] Lima, E. L. - Análise no Espaço ℝ^{n}. Brasília, Ed. Universidade de Brasília. São Paulo, Ed. E. Blücher Ltda., 1970.

[Lm2] Lima, E. L. - Curso de Análise (vol. II). Rio de Janeiro, Instituto de Matemática Pura e Aplicada, CNPq, 1981 (projeto Euclides).

[Lm3] Lima, E. L. - Espaços Métricos. Rio de Janeiro, Instituto de Matemática Pura e Aplicada, CNPq, 1977 (projeto Euclides).

[Lm4] Lima, E. L. - Elementos de Topologia Geral. Ao livro técnico S. A., 1970.

[LM] Lee, E. B. & Markus, L. - Foundations of Optimal Control Theory, Wiley, New York, 1968.

[LS] Liu, W. & Sussmann, H. J. - Shortest Paths for Sub-Riemannian Metrics on Rank-2 Distributions. Memoirs of the AMS. Number 564, Vol. 118, November, 1995.

[Mg] Montgomery, R. - A Survey of Singular Curves in Sub-Riemannian Geometry. J. Dyn. and Control Syst., 1:1:49-90, 1995.

[Pa] Palais, R. S. - Foundations of Global Non-Linear Analysis. W. A. Benjamin, Inc., New York, 1968.

[Pg] Pontryagin, L. S. - The Mathematical Theory of Optimal Processes. Gordon and Breach Science Publishers, 1986.

[Ru] Rudin, W. - Real and Complex Analysis, McGraw-Hill Book Company, San Francisco, 3\underline{a} ed., 1987.